A sample of pure CaCO3 weighing 0.2428g is dissolved in HCl and the solution diluted to 250ml in a volumetric flask. A 50.00ml aliquot requires 42.74ml of an EDTA solution for titration. A 200ml sample of water containing Ca+2 is titrated with 16.38ml of the EDTA solution. Calculate the degree of hardness of water in ppm CaCO3. 


1
Expert's answer
2021-01-27T04:36:12-0500

Solution:

1)

The reaction between CaCO3 and hydrochloric acid can be described using the following equation:

CaCO3(s) + 2HCl(aq) → Ca2+(aq) + 2Cl-(aq) + H2O(l) + CO2(g)

According to the equation: Moles of CaCO3 = Moles of Ca2+


Find the Moles of Ca2+:

(0.2428 g CaCO3) × (1 mol CaCO3/100.09 g CaCO3) × (1 mol Ca2+/1 mol CaCO3) = 0.002426 mol Ca2+

Moles of Ca2+ = 0.002426 mol Ca2+

Find the Molarity of Ca2+:

Molarity of Ca2+ = Moles of Ca2+ / Volume solution = 0.002426 mol / 0.250 L = 0.0097 M Ca2+ions

Molarity of Ca2+ = 0.0097 M Ca2+ ions


2)

When calcium ions (Ca2+) are titrated with EDTA a relatively stable calcium complex is formed:

Ca2+ + H2Y2- = CaY2- + 2H+

According to the equation: Moles of Ca2+ = Moles of H2Y2- (EDTA)


Find the Moles of EDTA:

Moles of EDTA = (0.0097 mol Ca2+ / L) × (0.050 L) × (1 mol H2Y2- / 1 mol Ca2+) = 0.000485 mol EDTA

Moles of EDTA = 0.000485 mol EDTA

Find the molarity of EDTA:

Molarity of EDTA = Moles of EDTA / Volume of EDTA = 0.000485 mol / 0.04274 L = 0.01135 M EDTA

Molarity of EDTA = 0.01135 M


3)

The hardness of water is calculated in terms of the weight of CaCO3 (1 ppm = 1 mg/L).


When calcium ions (Ca2+) are titrated with EDTA a relatively stable calcium complex is formed:

Ca2+ + H2Y2- = CaY2- + 2H+

According to the equation: Moles of Ca2+ = Moles of H2Y2- (EDTA)


Find the Moles of Ca2+:

Moles of Ca2+ = (0.01135 mol EDTA/L) × (0.01638 L/1) × (1 mol Ca2+/1 mol EDTA) = 0.000186 mol Ca2+

Moles of Ca2+ = 0.000186 mol Ca2+

Find the Mass of CaCO3:

Moles of CaCO3 = Moles of Ca2+ = 0.000186 mol

(0.000186 mol Ca2+) × (1 mol CaCO3/1 mol Ca2+) × (100.09 g CaCO3/1 mol CaCO3) = 0.01862 g CaCO3

Mass of CaCO3 = 0.01862 g

Find the ppm concentration CaCO3:

ppm CaCO3 = mg CaCO3 / L solution

ppm CaCO3 = (0.01862 g CaCO3) × (1000 mg CaCO3/1g CaCO3) × (1/0.200 L solution) = 93.1

ppm CaCO3 = 93.1


Answer: the degree of hardness of water is 93.1 ppm CaCO3